OpenStudy (anonymous):

(x^2-6x+9)/(x^2-9) ANS: (x-3)(x+3) Is this correct??

OpenStudy (anonymous):

100 per cent correct.

OpenStudy (mimi_x3):

(x-3) / (x+3)

OpenStudy (anonymous):

Thanks a lot.

OpenStudy (anonymous):

Thanks a lot, Diya and Mimi.

OpenStudy (anonymous):

As Mimi_x3 said. Factorise numerator and denominator, simplify. \[\Large \begin{array}{l} \frac{{{x^2} - 6x + 9}}{{{x^2} - 9}}\\ = \frac{{(x - 3)(x - 3)}}{{(x - 3)(x + 3)}}\\ = \frac{{(x - 3)}}{{(x + 3)}} \end{array}\]